OpenStudy (anonymous):

Evaluate. 5!/3! 20 60 120

OpenStudy (lgbasallote):

we can do this the long way... \[\frac{(5)(4)(3)(2)(1)}{(3)(2)(1)}\] cancel all that can be cancelled

OpenStudy (lgbasallote):

Hint: cancel the numbers that appear in your numerator AND denominator

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

5!/3! = 5x4x3x2x1/3x2x1 =120/6 =20... thts right.. its the 1st option..